Grade 12Algebra

Give first the step you will use to separate the variable and then solve the equation:
y – 4 = – 7

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

We have to add 4 to both the side of given equation,
Then we get,
= y - 4 +4 = - 7 + 4
= y = - 3

Thanks